# Picogram-Level Nanoplastic Analysis with Nanoelectromechanical System Fourier Transform Infrared Spectroscopy: NEMS-FTIR

### v2

@@ METHODS
As shown in Figure 1A, NEMS chips, or resonators, play a dual role as both sensing devices and sample holders, making them the core of NEMS–FTIR technology. The wafers used in this study to fabricate NEMS chips are made from silicon (Si) and SiN, with thicknesses of 380 m and 50 nm, respectively, and an instrinsic tensile stress of approximately 50 MPa for the SiN. After the final manufacturing step, the KOH etching, a 1x1 mm2 SiN membrane is released. This membrane has a perforated area with a diameter of approximately 600 m, consisting of 6 m holes spaced 3 m apart. It acts as the platform on which the sample is deposited for analysis. Two gold electrodes, each 10 m wide, are positioned across the chip to enable electrical transduction, allowing mechanical vibrations to be converted into electrical signals, as demonstrated by Bešić et al. [42] and Luhmann et al. [35].
@@
Nanoliter droplet dispensing
@@
A PIPEJET nanoDispenser (BioFluidix-Hamilton Freiburg, Germany), shown in Figure S7A, was used to dispense the dispersions on the membrane of the NEMS chips. Polyimide capillaries with an inner diameter of 200 m (Prod. No. PJ-20010) were used to deliver individual droplets of 20 nL. A 1 cm length of Tygon tubing (Prod. No. VM-20053-1) was attached to the capillary to facilitate sample loading. Liquid samples were inserted into the tubing using a PIPETMAN P1000 micropipette with matching DIAMOND D1000 TIPACK tips (Prod. No. F144059M and F171500, respectively, Gilson Incorporated, WI, USA). The stroke velocity was set to 90 m/ms during droplet dispensing.
For each concentration of the dilution series, droplets of 20 nL were deposited onto the center of the individual NEMS chips. The total deposited mass of each sample is listed in Table 2B. Each deposition was performed in triplicate, with three NEMS chips prepared for each specific polymer mass. Nine additional NEMS chips were sampled with blank samples to provide a reliable baseline for subtraction.
For the mixture of nanoplastic particles, a single replicate was prepared by depositing one 20 nL drop of the pre–mixed dispersion onto a NEMS chip. Blank samples were deposited in parallel onto three additional NEMS chips.
For the analysis of the tea bag leachates, three NEMS chips were sampled with leachate samples and three with blank samples using the PIPEJET nanoDispenser. Each NEMS chip was sequentially loaded with ten drops (10 nL each) of liquid sample, with a two–minute delay between drops, totaling 100 nL per chip.
Furthermore, three NEMS chips were prepared using the Drop casting Accessory (Invisible–Light Labs GmbH, Austria), a PIPETMAN P2 micropipette, and corresponding DIAMOND DL10 TIPACK tips (Prod. No. F144054M and F171200, respectively, Gilson Incorporated, WI, USA). A single 500 nL droplet of the tea bag leachate was deposited onto each chip. The chips were then left to dry inside the Drop casting Accessory, shown in Figure S7B, in an environment with a controlled humidity gradient, which, due to the presence of perforations in the chip, facilitates droplet drying within the perforated area. In parallel, three additional NEMS chips were sampled using blank samples.
@@
A system incorporating the Portable Aerosol Generation System (PAGS, Handix Scientific Inc., CO, USA) was employed to sample particles released from the nylon tea bags subjected to accelerated weathering cycles. This setup enabled consistent aerosolization and deposition of particles onto the NEMS chips facilitated by the holes in the central region of the membrane. Sampling was conducted at regular intervals (day 1, day 2, day 5, day 10, day 15, day 20, and day 25) to monitor the temporal evolution of particle release. The NEMS chip was placed in a holder (Aerosol Flow Adapter, Invisible–Light Labs GmbH, Austria), which was then positioned in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA) equipped with a built-in pump. The pump was set to a flow rate of 0.5 L/min for 5 minutes for each NEMS chip, which allowed sample deposition onto the membranes through the inertial impaction mechanism [38]. A schematic of the setup is presented in Figure S7C.
@@
A Hitachi SU8030 SEM (Hitachi, Japan) was employed to confirm the sizes of the particles deposited on the NEMS chips and to analyze the leachates from the plastic tea bags. For imaging, the sampled NEMS chips were mounted on an aluminum sample holder. Secondary electron (SE) detection in the upper lens mode was utilized to capture high-resolution surface details. The microscope operated at an accelerating voltage of 2 kV with an emission current of 3900 nA. NEMS chips were examined at various magnifications to assess the particle dimensions and the structural characteristics of the deposited material.
NEMS–FTIR spectroscopy
A nanoelectromechanical IR analyzer (EMILIE, Invisible–Light Labs GmbH, Austria) in conjunction with an FTIR spectrometer (Vertex 70, Bruker Optics, MA, USA), as shown in Figure 1B, was used to collect all the NEMS–FTIR spectra and confirm the chemical identity of different types of nanoplastic particles, their mixture, and tea bag leachates.
During the measurement process, sampled NEMS chips were placed in the vacuum chamber of the nanoelectromechanical IR analyzer at 10-5 mbar and the temperature of the chips was consistently regulated to 25C using the integrated thermoelectric cooling (TEC) controller and Peltier element. To minimize atmospheric interferences, the optical path of the EMILIE and FTIR spectrometer were purged with dry nitrogen throughout the measurements. A resolution of 4 cm-1, with a stabilization delay of 30 ms, 200 co-additions, and an aperture of 6 mm, were used for the FTIR parameters. The spectral range extended from 4000 to 400 cm-1.
ATR–FTIR spectroscopy
ATR–FTIR spectroscopy was used alongside the NEMS–FTIR to analyze the tea bag leachates as a reference IR technique. The tea bag material itself was also examined using ATR-FTIR to confirm it was the source of the detected components in the leachates. A Platinum ATR accessory with a diamond crystal (A225/Q, Bruker Optics, MA, USA) was used for all ATR measurements, performed on the same Bruker Vertex 70 spectrometer utilized for the NEMS–FTIR experiments.
To match the NEMS–FTIR experimental conditions, an aliquot of 500 nL of tea bag leachate was applied to the isopropanol–cleaned ATR diamond crystal. The droplet was dispensed using the same micropipette and tip combination as that used for sampling the NEMS chips with the same tea bag leachates (as described in the Nanoliter droplet dispensing section). Once the droplet had evaporated, the ATR measurement was performed by lowering the pressure clamp onto the remaining residue. Additionally, 500 nL of the corresponding blank sample was measured in the same manner.
@@
The ATR–FTIR spectra were processed using the "Extended ATR correction" (standard mode) and "Baseline correction" features in the OPUS software to correct for the slight distortions in peak intensity, shape, and position inherent to ATR-FTIR measurements [60, 61, 62, 63, 64]. The spectra were also smoothed using Savitzky-–Golay filtering (window length 20, polyorder 2).
@@
All raw NEMS–FTIR spectra were smoothed using a Savitzky–Golay filter with a window length of 20 and a polynomial order of 2. The spectra of the blanks and analytes were then divided by the recorded NEMS–FTIR spectrum of the FTIR light source to account for the wavelength-dependent intensity of the light source. The NEMS–FTIR spectrum of the light source was recorded using a NEMS chip coated with a 5 nm ultra-thin Pt film (EMILIE LIGHT chip, Invisible–Light Labs GmbH, Austria), which acts as a broadband absorber [77].
@@
Spectral calibration was carried out using the intrinsic SiN absorption peak at 835 cm-1 as a reference. This calibration was based on transmission measurements through 12 clean, perforated, mm2, 50 nm thick low–stress SiN NEMS membranes. Figure S8 in the Supporting Information shows the average transmittance and corresponding absorptance of the NEMS chip’s SiN material measured using the internal detector of the FTIR spectrometer (2 mm aperture, 32 scans, and a resolution of 2 cm-1). Despite the different sizes of the membranes ( mm2 and mm2) and perforation areas (600 m and 1200 m, respectively), the ratio of perforated to non–perforated membrane material remains consistent across both membrane types, making the calibration based on the SiN absorption peak possible.
@@
The absorptance for a specific wavenumber , , was calculated from the transmittance as , assuming insignificant scattering, yielding . The NEMS–FTIR response of the bare SiN chip, , was then related to absorptance through a calibration factor :
@@
Depositing the nanoplastics on the membrane with the nanodropper created a circular sample spot (see Figure S1 in Supporting Information). The nanoplastics occupied the area , which is smaller than the area of the IR beam illuminating the chip, . Furthermore, the increased photothermal response, if the sample is concentrated in the membrane center instead of being evenly distributed over the entire membrane, was corrected by [47]. The sample absorptance can now be calculated from the measured NEMS-FTIR signal of the sample, , as:
can be calculated from the reflective focal length ratio of the FTIR spectrometer and EMILIE optics, which here was . The resulting IR beam diameter on the NEMS chip is then given by . For the aperture of mm, used in all experiments, the IR beam diameter on the chip is approximately mm. The total area illuminated by the IR beam can be calculated as , with the area of the perforation .
@@
corresponds to the diameter of the central nanoplastics area on the membrane with a diameter of m, based on six independent measurements. The total sample area is then given by , with the fill factor of the perforation .
@@
was calculated by finite element method (FEM) simulations by Kanellopulos et al. [78]. The results for a square low-stress SiN membrane with a side length of 1 mm are given in Figure S9 in the Supporting Information, yielding for the given sample spot size. The nanoplastics spots on the membrane formed a coffee ring, as seen in (see Figure S1 in the Supporting Information). Up to 90% of particles typically accumulate at the droplet perimeter in the coffee ring, as shown for nanometer-sized PS dispersions [79, 80]. FEM simulations have shown that the formation of a coffee ring has no significant effect on the NEMS–FTIR response.
@@
This framework aims to estimate the actual mass of the nanoplastic on the NEMS chip responsible for the NEMS-FTIR signal. Following the approach of Dudani and Takahama [81], it was assumed that particle scattering is negligible compared to absorption; thus, correction factors related to scattering are not required. Particles are much smaller than the wavelength of IR radiation used in this study, placing them in the Rayleigh regime, where absorption dominates IR attenuation. In this case, the absorbance of the deposited particles can be expressed as a function of the linear attenuation coefficient :
@@
In NEMS–FTIR, the measured absorptance values are small and the absorbance (3) can be approximated using a first-order Taylor series expansion:
@@
where is the NEMS–FTIR response of the averaged blank spectra, and denotes the wavenumber corresponding to a characteristic vibrational mode of the sample.
@@
A detailed derivation of Eq. (7) can be found in the Supporting Information, section C.
@@
Lists of characteristic peak positions (in wavenumbers) related to nanoplastic particles and tea bag leachates, derivation of the decadic linear attenuation coefficient for sparse particle film, SEM images of the coffee ring effect, NEMS-FTIR spectra of polypropylene and polyvinyl chloride nanoparticles, SEM images of micro- and nanofragments from tea bag leachates, comparison of non-corrected and corrected ATR-FTIR spectrum of nylon tea bag material, comparison between NEMS-FTIR and ATR-FTIR spectra regarding the spectral artifacts, experimental setup for the accelerated aging of the tea bags, sampling setups, IR spectral features of silicon nitride membranes, dependence of relative photothermal responsivity on the IR beam radius, values for deposited and estimated mass of polystyrene nanoparticles along with the standard deviations

### v3

@@ METHODS
As shown in Fig. 1B, NEMS chips, or resonators, play a dual role as both sensing devices and sample holders, making them the core of NEMS-FTIR technology. The wafers used in this study to fabricate NEMS chips are made from silicon (Si) and SiN, with thicknesses of 380 m and 50 nm, respectively, and an intrinsic tensile stress of approximately 50 MPa for the SiN. After the final manufacturing step, the KOH etching, a 1x1 mm2 SiN membrane is released. This membrane has a perforated area with a diameter of approximately 600 m, consisting of 6 m holes spaced 3 m apart. It acts as the platform on which the sample is deposited for analysis. Two gold electrodes, each 10 m wide, are positioned across the chip to enable electrical transduction, allowing mechanical vibrations to be converted into electrical signals, as demonstrated by Bešić et al. [47] and Luhmann et al. [40] After fabrication, the disposable NEMS chips are stored in individual closed plastic containers to prevent cross-contamination and non-specific adsorption from the environment.
@@
Drop casting methods
@@
In the nanodroplet dispensing approach, a piezoelectric nanodroplet dispenser (PIPEJET nanoDispenser, BioFluidix-Hamilton Freiburg, Germany), shown in Fig. S12A, was used to deposit the dispersions on the membrane of the NEMS chips. Polyimide capillaries with an inner diameter of 200 m (Prod. No. PJ-20010) were used to deliver individual droplets of 20 nL. A 1 cm length of Tygon tubing (Prod. No. VM-20053-1) was attached to the capillary to facilitate sample loading. Liquid samples were inserted into the tubing using a micropipette (PIPETMAN P1000, Prod. No. F144059M, Gilson Incorporated, WI, USA) with matching micropipette tips (DIAMOND D1000 TIPACK tips, Prod. No. F171500, Gilson Incorporated, WI, USA). The stroke velocity was set to 90 m/ms during nanodroplet dispensing. For each concentration of the dilution series, droplets of 20 nL were deposited onto the center of the individual NEMS chips. The total deposited mass of each sample is listed in Table 3B. Each deposition was performed in triplicate, with three NEMS chips prepared for each specific polymer mass. Nine additional NEMS chips were sampled with blank samples to provide a reliable baseline for subtraction. For the mixture of nanoplastic particles, a single replicate was prepared by depositing one 20 nL drop of the pre-mixed dispersion onto a NEMS chip. Blank samples were deposited in parallel onto three additional NEMS chips. For the analysis of the teabag leachates in water, three NEMS chips were sampled with leachate samples and three with blank samples using the piezoelectric nanodroplet dispenser. Each NEMS chip was sequentially loaded with ten drops (10 nL each) of liquid sample, with a two-minute delay between drops, totaling 100 nL per chip.
In the approach combining the drop casting and the pervaporation method, three NEMS chips were prepared using the Drop Casting Accessory (Invisible-Light Labs GmbH, Austria), a micropipette (PIPETMAN P2, Prod. No. F144054M, Gilson Incorporated, WI, USA), and corresponding micropipette tips (DIAMOND DL10 TIPACK, Prod. No. F171200, Gilson Incorporated, WI, USA). A single 500 nL droplet of the teabag leachate was deposited onto each chip. The chips were then left to dry inside the Drop casting Accessory, as shown in Fig. S12B, in an environment with a controlled humidity gradient, which, due to the presence of perforations in the chip, facilitates droplet drying within the perforated area via pervaporation. The approximate drying time for a 500 nL droplet under these conditions was 30 minutes. In parallel, three additional NEMS chips were sampled using blank samples.
For the analysis of the complex aqueous matrices, single replicate samples were prepared using the piezoelectric nanodroplet dispenser by depositing three drops of 10 nL each, totaling 30 nL, onto each NEMS chip. The liquid samples consisted of unprocessed tea leaves and teabag extracts prior to any sample preparation steps as well as after organic digestion and ultrafiltration. Corresponding single replicate blank samples, prepared from tea leaves extracts without teabag material, were deposited in the same manner.
@@
A system incorporating the Portable Aerosol Generation System (PAGS, Handix Scientific Inc., CO, USA) was employed to sample particles released from the nylon teabags subjected to accelerated weathering cycles. This setup enabled consistent aerosolization and deposition of particles onto the NEMS chips facilitated by the holes in the central region of the membrane. Sampling was conducted at regular intervals (day 1, day 2, day 5, day 10, day 15, day 20, and day 25) to monitor the temporal evolution of particle release. The NEMS chip was placed in a holder (Aerosol Flow Adapter, Invisible-Light Labs GmbH, Austria), which was then positioned in an eight-channel filter sampler (FILT, Brechtel Manufacturing Inc., CA, USA) equipped with a built-in pump. A schematic of the setup is presented in Fig. S12C. After nebulization, the aerosolized droplets containing the extracted particles are drawn through the central perforation of the NEMS chip. These small openings constrain and accelerate the airflow, increasing particle momentum. Particles with sufficient inertia no longer follow the bending streamlines around the membrane but instead depart from the flow and impact directly onto the perforated area.[43, 41, 40] Due to the inertial impaction mechanism, the deposited sample covers the entire perforated area of the membrane, as shown in Fig. S1B. The pump was set to a flow rate of 0.5 L/min for 5 minutes for each NEMS chip.
@@
SEM (Hitachi SU8030, Hitachi, Japan) was employed to confirm the sizes of the particles deposited on the NEMS chips and to analyze the leachates from the plastic teabags. For imaging, the sampled NEMS chips were mounted on an aluminum sample holder. Secondary electron (SE) detection in the upper lens mode was utilized to capture high-resolution surface details. The microscope operated at an accelerating voltage of 2 kV with an emission current of 3900 nA. NEMS chips were examined at various magnifications to assess the particle dimensions and the structural characteristics of the deposited material.
NEMS-FTIR spectroscopy
A nanoelectromechanical IR analyzer (EMILIE, Invisible-Light Labs GmbH, Austria) in conjunction with an FTIR spectrometer (Vertex 70, Bruker Optics, MA, USA), as shown in Fig. 1A, was used to collect all the NEMS-FTIR spectra and confirm the chemical identity of different types of nanoplastic particles, their mixture, and teabag leachates.
During the measurement process, sampled NEMS chips were placed in the vacuum chamber of the nanoelectromechanical IR analyzer at 10-5 mbar and the temperature of the chips was consistently regulated to 25C using the integrated thermoelectric cooling (TEC) controller and Peltier element. To minimize atmospheric interferences, the optical path of the EMILIE and FTIR spectrometer were purged with dry nitrogen throughout the measurements. A resolution of 4 cm-1, with a stabilization delay of 30 ms, 200 co-additions, and an aperture of 6 mm, were used for the FTIR parameters. The spectral range extended from 4000 to 400 cm-1. The spectrometer aperture was chosen to be sufficiently large, such that the IR beam spot is always significantly larger than the membrane area. For each individual chip, the IR beam was fine-aligned to ensure maximal signal amplitude.
ATR-FTIR spectroscopy
ATR-FTIR spectroscopy was used alongside the NEMS-FTIR to analyze the teabag leachates as a reference IR technique. The teabag material itself was also examined using ATR-FTIR to confirm it was the source of the detected components in the leachates. A Platinum ATR accessory with a diamond crystal (A225/Q, Bruker Optics, MA, USA) was used for all ATR measurements, performed on the same Bruker Vertex 70 spectrometer utilized for the NEMS-FTIR experiments. All ATR-FTIR measurements were performed under nitrogen-purged conditions to ensure consistency with the NEMS-FTIR measurement environment.
To match the NEMS-FTIR experimental conditions, an aliquot of 500 nL of teabag leachate was applied to the isopropanol-cleaned ATR diamond crystal. The droplet was dispensed using the same micropipette and tip combination as that used for sampling the NEMS chips with the same teabag leachates (as described in the "Drop casting methods" section). Once the droplet had evaporated, the ATR measurement was performed by lowering the pressure clamp onto the remaining residue. Additionally, 500 nL of the corresponding blank sample was measured in the same manner.
@@
The ATR-FTIR spectra were processed using the "Extended ATR correction" (standard mode) and "Baseline correction" features in the OPUS software to correct for the slight distortions in peak intensity, shape, and position inherent to ATR-FTIR measurements. [54, 55, 56, 57, 58] The spectra were also smoothed using Savitzky–Golay filtering (window length 20, polyorder 2).
@@
All raw NEMS-FTIR spectra were smoothed using a Savitzky-Golay filter with a window length of 20 and a polynomial order of 2. The spectra of the blanks and analytes were then divided by the recorded NEMS-FTIR spectrum of the FTIR light source to account for the wavelength-dependent intensity of the light source. The NEMS-FTIR spectrum of the light source was recorded using a NEMS chip coated with a 5 nm ultra-thin Pt film (EMILIE LIGHT chip, Invisible-Light Labs GmbH, Austria), which acts as a broadband absorber. [106] Following this correction, the spectra were scaled by a normalization factor to set the SiN peak at 835 cm-1 to a value of 1 .
@@
Spectral calibration was carried out using the intrinsic SiN absorption peak at 835 cm-1 as a reference. This calibration was based on transmission measurements through 12 clean, perforated, mm2, 50 nm thick low-stress SiN NEMS membranes. Fig. S4 in the Supporting Information shows the average transmittance and corresponding absorptance of the NEMS chip’s SiN material measured using the internal detector of the FTIR spectrometer (2 mm aperture, 32 scans, and a resolution of 2 cm-1). Despite the different sizes of the membranes ( mm2 and mm2) and perforation areas (600 m and 1200 m, respectively), the ratio of perforated to non-perforated membrane material remains consistent across both membrane types, making the calibration based on the SiN absorption peak possible.
@@
The absorptance for a specific wavenumber , , was calculated from the transmittance as , assuming insignificant scattering, yielding . The NEMS-FTIR response of the bare SiN chip, , was then related to absorptance through a calibration factor :
@@
Depositing the nanoplastics on the membrane using the piezoelectric nanodroplet dispenser created a circular sample spot (see Fig. S2 in Supporting Information). The nanoplastics occupied the area , which is smaller than the area of the IR beam illuminating the chip, . Furthermore, the increased photothermal response, if the sample is concentrated in the membrane center instead of being evenly distributed over the entire membrane, was corrected by . [52] The sample absorptance was calculated from the measured NEMS-FTIR signal of the sample, , as:
was calculated based on the experimentally determined effective IR beam diameter, . For each deposited mass of the PS nanoparticles, was iteratively adjusted to optimize the fit between the measured NEMS-FTIR spectrum and a reference spectrum generated from the known refractive index of PS. [73] The optimization focused particularly on matching the intensity and shape of the characteristic absorption peak at 1452 cm-1. This method provided an estimate of the effective beam diameter, which represents the actual size of the IR spot that effectively illuminates the SiN membrane. was found to be mm, which is close to the ideal value corresponding to the SiN membrane with a lateral dimension of 1 mm, and the total area illuminated by the IR beam was calculated as , with the area of the perforation .
@@
corresponds to the diameter of the central nanoplastics area on the membrane with a diameter of m. These measurements were obtained by analyzing six separate NEMS chips, each imaged under a microscope. was then quantified from these images using ImageJ software. The total sample area is then given by , with the fill factor of the perforation .
@@
was calculated by finite element method (FEM) simulations by Kanellopulos et al. [107] The results for a square low-stress SiN membrane with a side length of 1 mm are given in Fig. S13 in the Supporting Information, yielding for the given sample spot size. Thus, the response is 1.68 times greater than it would be for a sample uniformly distributed across the entire square membrane. The nanoplastics spots on the membrane formed a coffee ring, as shown in Fig. S2 in the Supporting Information. Up to 90% of particles typically accumulate at the droplet perimeter in the coffee ring, as shown for nanometer-sized PS dispersions. [108, 109] FEM simulations have shown that the formation of a coffee ring has no significant effect on the NEMS-FTIR response (see Fig. S14), with an estimated uncertainty in the responsivity of < 3%.
@@
For samples containing a complex matrix, a modified subtraction procedure was required to account for differences in the tea-leaf background between the blank and analyte spectra. Although both spectra were prepared from the same initial tea infusion by splitting it into two aliquots, where one was used as the blank and the other was enriched with an empty nylon teabag, the background contributions from the leaf extract were not identical. This likely stems from the heterogeneous nature of the organic material. Despite using fresh leaves to minimize fragmentation, small particulate or colloidal components may have been unevenly distributed between the aliquots. To eliminate the matrix-dependent spectral offset, spectra were normalized over a region between 3700 and 2500 cm-1 and subtracted. As a result, the processed spectra are presented in arbitrary units.
@@
This framework aims to estimate the actual mass of the nanoplastic on the NEMS chip responsible for the NEMS-FTIR signal. Following the approach of Dudani and Takahama, [110] it was assumed that particle scattering is negligible compared to absorption; thus, correction factors related to scattering are not required. Particles are much smaller than the wavelength of IR radiation used in this study, placing them in the Rayleigh regime, where absorption dominates IR attenuation. In this case, the absorbance of the deposited particles can be expressed as a function of the linear attenuation coefficient :
@@
In NEMS-FTIR, the measured absorptance values are small and the absorbance (3) can be approximated using a first-order Taylor series expansion:
@@
where is the NEMS-FTIR response of the averaged blank spectra, and denotes the wavenumber corresponding to a characteristic vibrational mode of the sample.
@@
A detailed derivation of Eq. (7) can be found in the Supporting Information, section A.
@@
Derivation of the decadic linear attenuation coefficient for sparse particle film, sample distribution on the membrane of the NEMS chip after aerosol-based deposition and drop casting in conjunction with pervaporation, sample distribution on the membrane of NEMS chip after using the nanodroplet dispenser to deposit the nanoparticles, SEM images of polypropylene and polyvinyl chloride nanoparticles sitting on the perforated area of the membrane, IR spectral features of silicon nitride membranes, NEMS-FTIR spectra of polypropylene nanoparticles, NEMS-FTIR spectra of polyvinyl chloride nanoparticles, correlation between the deposited (amount of deposited PS particles) and estimated (from the measured absorptances of PS particles) mass of nanoplastic particles, SEM images of micro- and nanofragments from teabag leachates, close-up view of NEMS-FTIR and ATR-FTIR spectra of a 500 nL teabag leachate, comparison of non-corrected and corrected ATR-FTIR spectrum of bulk nylon teabag, comparison between NEMS-FTIR and ATR-FTIR spectra regarding the spectral artifacts, experimental setup for the accelerated aging of the teabags, sampling setups, dependence of relative photothermal responsivity on the IR beam radius, FEM simulation presenting temperature distribution in a silicon nitride membrane induced by laser heating, table of characteristic peak positions (in wavenumbers) related to nanoplastic particles and nylon teabag leachates, table of values for deposited and estimated mass of polystyrene nanoparticles along with the standard deviations.